OpenStudy (anonymous):

10. Simplify the rational expression by rationalizing the demominator. 3 sqrt 160/sqrt 1350x A) 2 sqrt 15x/15x B) 3 sqrt 160/1350x C) 4 sqrt 15x/15x D) 4/ sqrt 45x 11. Find the mean, median, and mode of the data set. Round to the nearest tenth. 12, 23, 17, 20, 14, 14, 18, 16, 17, 18, 18 A) mean= 15.6, median=17 mode=18 B) mean= 17, median=18, mode=17 C) mean= 15.6, median=18, mode=17 D) mean=17, median=17, mode=18

OpenStudy (boldjon):

i'm here i'm here

OpenStudy (boldjon):

just ask the calculator for number 11 http://www.wolframalpha.com/input/?i=mean+12%2C23%2C17%2C20%2C14%2C14%2C18%2C16%2C17%2C18%2C18+

OpenStudy (boldjon):

not gonna lie i idn't really get 10

OpenStudy (susanna):

\[\sqrt[3]{160}\div \sqrt{1350x}?\]

OpenStudy (boldjon):

I believe it's d

OpenStudy (boldjon):

11 that is

OpenStudy (susanna):

or\[3\sqrt{160}\div \sqrt{1350x}\]

OpenStudy (boldjon):

want me to explain?